### v2

# Minimal-backaction work statistics of coherent engines

Abstract
Determining the work statistics of quantum engines is challenging due to measurement backaction. We here show that a dynamic Bayesian network-based measurement scheme, which preserves quantum coherence within an engine cycle, is minimally invasive, in the sense that the averaged measured state over one cycle exactly coincides with the unmeasured state. It therefore provides a general framework to investigate energy exchange statistics in quantum machines. This stands in contrast to the standard two-point measurement protocol, whose backaction can be so strong that it generally fails to reproduce the average work output of a coherent motor. It may even alter its mode of operation, causing it to cease functioning as an engine under observation. We further demonstrate that recently proposed universal fluctuation bounds do not necessarily apply to coherent machines.
Microscopic machines are subject to thermal sch03 ; kay07 ; han09 ; kas17 and, at low enough temperatures, to additional quantum fluctuations kos14 ; ben17 ; mye22 ; can24 . These quantum fluctuations arise from random transitions between discrete energy levels during the machine’s evolution, and give rise to distinctly nonclassical features mukamel_2009 ; talkner_2011 . As a result, common thermodynamic figures of merit, such as power output and efficiency callen_1985 , become stochastic variables. Because the presence of fluctuations strongly influences the performance of small-scale engines vdbroeck_2014 ; esposito_2015 ; seifert_2018 ; ryabov_2018 ; krishnamurthy_2019 ; hol21 , from biological to synthetic motors sch03 ; kay07 ; han09 ; kas17 ; kos14 ; ben17 ; mye22 ; can24 , a detailed characterization of their random properties is essential. However, whereas the fluctuation statistics of classical microscopic engines may be experimentally determined without affecting their operation bli12 ; mar15 ; kri16 ; mar17 , this is generally not possible in the quantum regime, where measurement backaction inevitably perturbs the system jacobs_2014 . A central issue is therefore to probe the fluctuations of a quantum machine with minimal disturbance.
Quantum coherence, encoded in the off-diagonal density matrix elements, is a fundamental attribute of nonclassical thermodynamic transformations plenio_2017 . It is, however, a fragile resource that can be readily destroyed by external perturbations, making the investigation of energy fluctuations in coherent machines inherently difficult acin_2017 ; wu19 ; sampaio_2019 ; paternostro_2020 ; trombettoni_2021 ; imparato_2024 ; skrzypczyk_2025 . The standard two-point measurement scheme mukamel_2009 ; talkner_2011 used to evaluate the quantum statistics of energy exchanges relies on projective energy measurements performed at the beginning and at the end of a process hanggi_2007 . Owing to its projective nature, and the associated measurement backaction, this scheme fails to properly account for quantum coherence in the energy basis. It is thus limited to the analysis of fluctuations in incoherent motors den20 ; jia21 ; den21 ; lutz_2021 ; fei22 ; serra_2024 . A prominent alternative is the framework of dynamic Bayesian networks which provides a general method to analyze conditional dependencies among time-dependent random quantities, from engineering and robotics to finance and biological networks neapolitan_2004 ; darwiche_2009 . In the quantum setting, this approach enables the inference of energy fluctuations from indirect measurements through conditional probabilities evaluated using Bayes’ rule lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 . In contrast to the two-point measurement protocol, Bayesian networks fully capture quantum coherence throughout the evolution lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 .
In this article, we investigate the energy exchange statistics in a generalized finite-time quantum Otto qudit engine in which coherence is dynamically generated along the cycle kos17 . This quantum machine can be viewed as a multidimensional extension of recently realized qubit-based engines ass19 ; pet19 ; hou25 . We establish that the Bayesian network approach is minimally invasive, in the sense that the average state of the engine after a round of measurements is identical to the unmeasured state. The Bayesian network scheme therefore induces minimal measurement backaction, in sharp contrast to the two-point measurement technique. We further show that the discrepancy between the two methods is directly linked to the presence of quantum coherence in the state of the working substance. Remarkably, the backaction associated with the two-point measurement scheme can be so strong that it qualitatively alters the operation regime of the otherwise unobserved quantum machine, driving a transition from engine to heater or accelerator. Finally, we demonstrate that recently proposed ”universal” bounds on engine fluctuations, derived for quasistatic Otto cycles agarwalla_2021_2 ; agarwalla_2021 ; das23 ; moh23 ; moh23a ; watanabe_2025 , can be violated in the presence of coherence.
Generalized coherent Otto cycle.
We consider a quantum spin engine whose working substance $S$ is described by the density operator $\rho_{i}$ and a time-dependent Hamiltonian $H_{i}(t)=\omega_{i}(t)S_{z}+g_{i}(t)S_{x}$, where $S_{x,y,z}$ are the spatial components of the $d_{S}$-dimensional spin operator, and the index $i$ labels the four branches of the cycle. The function $\omega_{i}(t)$ controls the energy spacing of $H_{i}$, while $g_{i}(t)$ modulates the amount of coherence in the basis of $S_{z}$, since $S_{z},S_{x}]\neq 0$. The engine performs a finite-time quantum Otto cycle during which the working medium is repeatedly and weakly coupled davies_1974 to a hot (cold) heat reservoir at temperature $T_{\mathrm{h}}$ ($T_{\mathrm{c}}$). To make system independent statements, we characterize the bath interaction with a generalized amplitude damping map $\mathcal{T}_{t}$ that leads to full thermalization in the infinite-time limit nie10 ; srikanth_2024 .
The four branches of the cycle are as follows kos17 : (1) Isentropic compression: the system $S$ undergoes unitary evolution for a time $t_{\mathrm{k}}$, from state $\rho_{1}$ to state $\rho_{2}=U_{\mathrm{k}}(t_{\mathrm{k}})\rho_{1}U_{\mathrm{k}}^{\dagger}(t_{\mathrm{k}})$, where $U_{\mathrm{k}}$ is the time evolution operator generated by $H_{\mathrm{k}}$. The control parameters $\omega_{\mathrm{k}}$ and $g_{\mathrm{k}}$ are chosen such that $\omega_{\mathrm{k}}(0)=\omega_{\mathrm{c}}$ and $\omega_{\mathrm{k}}(t_{\mathrm{k}})=\omega_{\mathrm{h}}$ ($\omega_{\mathrm{c}}$), and $g_{\mathrm{k}}(0)=g_{\mathrm{k}}(t_{\mathrm{k}})=0$. (2) Hot isochore: the working substance $S$ is coupled to the hot bath for a finite time $t_{\mathrm{h}}$, leading to the state $\rho_{3}=\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}(\rho_{2})$. In the infinite-time limit, the system asymptotically thermalizes to the Gibbs state $\mathcal{T}_{\infty}^{\mathrm{h}}(\rho_{2})=\text{exp}[-H_{\mathrm{k}}(t_{\mathrm{k}})/T_{\mathrm{h}}]/\text{tr}\{\text{exp}[-H_{\mathrm{k}}(t_{\mathrm{k}})/T_{\mathrm{h}}]\}$. (3) Isentropic expansion: during unitary evolution for time $t_{\mathrm{e}}$, the state $\rho_{3}$ transitions to state $\rho_{4}=U_{\mathrm{e}}(t_{\mathrm{e}})\rho_{3}U_{\mathrm{e}}^{\dagger}(t_{\mathrm{e}})$, where $U_{\mathrm{e}}$ is generated by the Hamiltonian $H_{\mathrm{e}}$. The parameters $\omega_{\mathrm{e}}$ and $g_{\mathrm{e}}$ here satisfy $\omega_{\mathrm{e}}(0)=\omega_{\mathrm{h}}$ and $\omega_{\mathrm{e}}(t_{\mathrm{e}})=\omega_{\mathrm{c}}$, as well as $g_{\mathrm{e}}(0)=g_{\mathrm{e}}(t_{\mathrm{e}})=0$. (4) Cold isochore: in the last stroke, the system interacts with the cold reservoir for a finite time $t_{\mathrm{c}}$, so that the final state is given by $\rho_{5}=\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}(\rho_{4})$. In this case, the asymptotic equilibrium state reads $\mathcal{T}_{\infty}^{\mathrm{c}}(\rho_{4})=\text{exp}[-H_{\mathrm{e}}(t_{\mathrm{e}})/T_{\mathrm{c}}]/\text{tr}\{\text{exp}[-H_{\mathrm{e}}(t_{\mathrm{e}})/T_{\mathrm{c}}]\}$.
One complete engine cycle is hence described by the map $\Lambda=\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}$, where $\mathcal{U}_{i}(\cdot)=U_{i}(t_{i})\cdot U_{i}^{\dagger}(t_{i})$ with a total duration of $\tau=t_{\mathrm{k}}+t_{\mathrm{h}}+t_{\mathrm{e}}+t_{\mathrm{c}}$. After an initial transient, the motor will reach a limit-cycle regime for the state $\rho_{1}$ characterized by the fixed point of the map, $\Lambda(\tilde{\rho}_{1})=\tilde{\rho}_{1}$ com . In general, this state will exhibit quantum coherence in the basis of $S_{z}$ for nonzero transverse drivings $g_{i}(t)$ and finite bath-coupling times $t_{i}$. Such coherence impacts the thermodynamic performance of the machine. The fluctuation analysis presented below will thus be based on these limit-cycle states.
We begin by recalling the thermodynamics of the unobserved, that is, unperturbed, quantum Otto engine kos17 . During the hot and cold isochoric strokes, the system Hamiltonian is held constant. As a consequence, no work is performed, and only heat is exchanged with the two thermal baths. In both cases, the exchanged heat is given by the variation of the mean (internal) energy of the working substance: $Q_{\mathrm{h}}=\omega_{\mathrm{h}}\text{tr}\{S_{z}[\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}(\tilde{\rho}_{1})-\mathcal{U}_{\mathrm{k}}(\tilde{\rho}_{1})]\}$ and $Q_{\mathrm{c}}=-\omega_{\textrm{c}}\text{tr}\{S_{z}[\mathcal{U}_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}(\tilde{\rho}_{1})-\tilde{\rho}_{1}]\}$. The total work then follows from the first law of thermodynamics as $W=Q_{\mathrm{h}}+Q_{\mathrm{c}}$ kos17 . Engine operation is characterized by a positive work output, $W>0$, together with the condition that heat is absorbed from the hot bath, $Q_{\textrm{h}}>0$, and released into the cold one, $Q_{\textrm{c}}<0$.
Fluctuation statistics of a coherent engine.
The distribution $P(w)$ of the stochastic work output of a quantum machine can be obtained by sequentially measuring the working substance at the four corners of the cycle. Its general form is den20 ; jia21 ; den21 ; lutz_2021 ; fei22 ; serra_2024
$P(w)$ $=\sum_{j,l,m,n}\delta[w+(e_{\mathrm{k}}^{l}-e_{\mathrm{e}}^{j})+(e_{\mathrm{e}}^{n}-e_{\mathrm{k}}^{m})]$ (1)
$\times p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n}),$
where $e_{\mathrm{k},\mathrm{e}}^{j}$ is the $j$-th eigenvalue of the Hamiltonian $H_{\mathrm{k},\mathrm{e}}(t_{\mathrm{k},\mathrm{e}})$ and $p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})$ denotes the joint probability of the four outcomes den20 ; jia21 ; den21 ; lutz_2021 ; fei22 ; serra_2024 . The latter quantity depends on the chosen measurement protocol, and hence on the associated measurement backaction.
To illustrate the difference between the two-point measurement (TPM) scheme mukamel_2009 ; talkner_2011 and the dynamic Bayesian network (DBN) approach neapolitan_2004 ; darwiche_2009 , we consider an arbitrary evolution taking a state $\rho_{\textrm{i}}$ of a system to another state $\rho_{\textrm{f}}=\mathrm{M}(\rho_{\textrm{i}})$, with $\mathrm{M}$ a CPTP map nie10 , while also changing its Hamiltonian from $H_{\textrm{i}}$ to $H_{\textrm{f}}$. The distribution of the random energy change $\Delta u$ of the system during this process is $P(\Delta u)=\sum_{j,k}\delta[\Delta u-(e_{\textrm{f}}^{k}-e_{\textrm{i}}^{j})]\,p(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$. Within the TPM scheme, the joint probability $p(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ is obtained by first performing a projective measurement in the eigenbasis of $H_{\textrm{i}}$ on $\rho_{\textrm{i}}$, and then measuring the evolved projected state in the eigenbasis of $H_{\textrm{f}}$ mukamel_2009 ; talkner_2011 : $p_{\textrm{TPM}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})=\text{tr}[\Pi_{\textrm{f}}^{k}\,\mathrm{M}(\Pi_{\textrm{i}}^{j}\rho_{\textrm{i}}\Pi_{\textrm{i}}^{j})]$, where $\Pi_{\textrm{i},\textrm{f}}^{j}$ is the eigenprojector corresponding to $e_{\textrm{i},\textrm{f}}^{j}$. The first measurement collapses $\rho_{\textrm{i}}$ into an eigenspace of $H_{\textrm{i}}$, and therefore destroys all its coherences in that basis. The same happens with the second projective measurement.
On the other hand, the dynamic Bayesian network method is based on projective measurements in the eigenbasis of $\rho_{\textrm{i},\textrm{f}}$, instead of that of $H_{\textrm{i},\textrm{f}}$, hence preserving initial and final coherences in the energy basis lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 . The joint probability $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ of the unmeasured energy eigenvalues is then obtained using Bayesian inference. Let us concretely write the eigendecomposition of the initial and final states as $\rho_{\textrm{i},\textrm{f}}=\sum_{\alpha}\lambda_{\textrm{i},\textrm{f}}^{\alpha}\mathrm{P}_{\textrm{i},\textrm{f}}^{\alpha}$ where $\mathrm{P}_{\textrm{i},\textrm{f}}^{\alpha}$ is the eigenprojector corresponding to the eigenvalue $\lambda_{\textrm{i},\textrm{f}}^{\alpha}$. By now performing the first projective measurement on the complete measurement ensemble, dividing it into two halves, time evolving one of them and then projectively measuring it, the joint distribution of measurement outcomes is $p(\lambda_{\mathrm{i}}^{\alpha},\lambda_{\mathrm{f}}^{\beta})=\text{tr}[\mathrm{P}_{\textrm{f}}^{\beta}\,\mathrm{M}(\mathrm{P}_{\textrm{i}}^{\alpha}\rho_{\textrm{i}}\mathrm{P}_{\textrm{i}}^{\alpha})]$. The next step is to evaluate the posterior conditional probabilities for each Hamiltonian eigenvalue using the standard Born rule, $p(e_{\mathrm{i}}^{j}|\lambda_{\mathrm{i}}^{\alpha})=\text{tr}[\Pi_{\textrm{i}}^{j}\mathrm{P}_{\textrm{i}}^{\alpha}\rho_{\textrm{i}}\mathrm{P}_{\textrm{i}}^{\alpha}/\text{tr}(\mathrm{P}_{\textrm{i}}^{\alpha}\rho_{\textrm{i}})]$ and $p(e_{\mathrm{f}}^{k}|\lambda_{\mathrm{f}}^{\beta})=\text{tr}[\Pi_{\textrm{f}}^{k}\mathrm{P}_{\textrm{f}}^{\beta}\rho_{\textrm{f}}\mathrm{P}_{\textrm{f}}^{\beta}/\text{tr}(P_{\textrm{f}}^{\beta}\rho_{\textrm{f}})]$. Summing over all possible measurement outcomes, one finds lutz_2020 ; par20 ; mic20a ; str20 ; zha22 ; lutz_2023 ; lutz_2024 , $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})=\sum_{\alpha,\beta}p(e_{\mathrm{i}}^{j}|\lambda_{\mathrm{i}}^{\alpha})\,p(e_{\mathrm{f}}^{k}|\lambda_{\mathrm{f}}^{\beta})\,p(\lambda_{\mathrm{i}}^{\alpha},\lambda_{\mathrm{f}}^{\beta})$. The two joint distributions, $p_{\textrm{TPM}}$ and $p_{\textrm{DBN}}$, coincide in the absence of quantum coherence in the energy basis, yielding identical energy statistics. However, because the states $\rho_{\textrm{i},\textrm{f}}$ are diagonal in their eigenbases, measurement backaction is strongly reduced in the DBN protocol in the presence of energy coherences, compared to the TPM approach. Importantly, the Bayesian network distribution $p_{\textrm{DBN}}(e_{\textrm{i}}^{j},e_{\textrm{f}}^{k})$ is accessible experimentally mic20a .
Figure 1: Entropic distance, $D_{\textrm{KL}}(P_{\textrm{DBN}}||P_{\textrm{TPM}})$, between the work distributions of the dynamic Bayesian network (DBN) approach and the two-point measurement (TPM) scheme versus the relative entropies of coherence, $\mathcal{C}_{1,3}$, after the two isochores in a three-level quantum Otto cycle. The largest difference is observed for maximal coherence. It vanishes for zero transverse driving strength $g=0$ (inset). Parameters are $\omega_{\mathrm{h}}=10$, $T_{\mathrm{h}}=14$, $\omega_{\mathrm{c}}=0.5$, $T_{\mathrm{c}}=0.1$, $g=9$, $\lambda_{\mathrm{c}}=\lambda_{\mathrm{h}}=\lambda$.
The joint probability $p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})$ in Eq. (1) can be derived for the quantum Otto engine in a similar manner by extending the previous considerations to the four corners of the cycle. The explicit (and lengthy) expressions for the work output distribution $P(w)$ for the two measurement schemes are given in the Supplemental Material. We here indicate the corresponding average work outputs $\langle w\rangle=\int dw\,w\,P(w)$. We have
$\langle w\rangle_{\mathrm{TPM}}$ $=$ $\omega_{\mathrm{h}}\text{tr}\{{J}_{z}[\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})-\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})]\}$ (2)
$-\omega_{\mathrm{c}}\text{tr}\{{J}_{z}[\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})-\tilde{\rho}_{1}]\},$
$\langle w\rangle_{\mathrm{DBN}}$ $=$ $\omega_{\mathrm{h}}\text{tr}\{{J}_{z}[\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}(\tilde{\rho}_{1})-\mathcal{U}_{\mathrm{k}}(\tilde{\rho}_{1})]\}$ (3)
$-\omega_{\mathrm{c}}\text{tr}\{{J}_{z}[\mathcal{U}_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}(\tilde{\rho}_{1})-\tilde{\rho}_{1}]\},$
where $\mathrm{D}_{\textrm{i}}(\cdot)=\sum_{j}\Pi_{\textrm{i}}^{j}\cdot\Pi_{\textrm{i}}^{j}$ is the dephasing map nie10 (for the TPM result, we have additionally used $\mathrm{D}_{\mathrm{k}},\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}]=0$).
Remarkably, the average work, $\langle w\rangle_{\mathrm{DBN}}$ in Eq. (3), obtained via Bayesian networks always agrees exactly with the unmeasured result $W$. By contrast, the two-point measurement expression, $\langle w\rangle_{\mathrm{TPM}}$ in Eq. (2), differs from it in general due to measurement backaction. A comparison of Eqs. (2) and (3) reveals that this discrepancy originates from the presence of quantum coherence during the cycle: $\langle w\rangle_{\mathrm{TPM}}=\langle w\rangle_{\mathrm{DBN}}=W$ only if the states of the working substance after the hot and cold isochores are diagonal in the eigenbasis of their respective Hamiltonians, that is, $\tilde{\rho}_{1}=\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})$ after the cold isochore and $\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})=\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})$ after the hot isochore. For the specific Hamiltonian considered here, this is achieved either in the absence of the transverse driving ($g_{i}=0$), or when the system fully thermalizes with the baths ($t_{\mathrm{h},\mathrm{c}}\to\infty$). In these cases the two work distributions (1) also coincide, $P_{\mathrm{TPM}}(w)=P_{\mathrm{DBN}}(w)$. For arbitrary system Hamiltonians $H_{i}$, and finite thermalization times, these conditions can be generically formulated as (Supplemental Material)
$\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\rho)\!=\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\rho)\text{ and }\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\rho)\!=\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\rho)\!\!$ (4)
since the output of the generalized amplitude damping map $\mathcal{T}_{t}$ is diagonal only if the input is already diagonal. Equation (4) is satisfied whenever the evolution operators $\mathcal{U}_{\mathrm{k,e}}$ do not induce coherences on incoherent states. It is worth noting that dynamic Bayesian networks are different from so-called collective measurements which have been shown to reduce measurement backaction acin_2017 ; wu19 . Whereas the measured average work in that scheme approaches the unmeasured value, after proper parameter optimization, it does not reach it in general.
[…]
Figure 2: Trace distance, $d(\langle\rho\rangle,\tilde{\rho}_{1})$, between the averaged measured state $\langle\rho\rangle$ and the unmeasured steady state $\tilde{\rho}_{1}$ for the two measurement protocols (TPM and DBN). It vanishes for dynamic Bayesian networks, showing that the latter minimizes measurement backaction. Same parameters as in Fig. 1.
Figure 3: Operation regimes of the quantum Otto machine for the two measurement schemes. a) Unmeasured and DBN-measured machines have identical mean work and heat, and hence display identical behavior. b) By contrast, the TPM-measured machine exhibits radically modified regimes due to measurement backaction. Same parameters as in Fig. 1.
Figure 4: The fluctuation bound, $\eta^{(2)}/\eta_{C}^{2}\leq 1$, Eq. (5), can be violated both for the TPM and DBN protocols owing to quantum coherence within the engine cycle for small thermalization rates $\lambda$, in the quasistatic regime. Parameters are $\omega_{\mathrm{h}}=1$, $T_{\mathrm{h}}=1.2$, $\omega_{\mathrm{c}}=0.85$, $T_{\mathrm{c}}=1$, $g=0.06$, and $\lambda_{\mathrm{c}}=\lambda_{\mathrm{h}}=\lambda$.
Minimal measurement backaction.
Both measurement protocols are unavoidably accompanied by measurement backaction. However, the above findings suggest that the backaction induced by the dynamic Bayesian network scheme is significantly weaker than that of the two-point measurement approach. To quantify the disturbance induced by a measurement sequence, we compare the unmeasured steady state $\tilde{\rho}_{1}$ with the averaged state obtained after a full engine cycle interspersed by five measurements, back to the initial corner of the cycle. We find $\langle\rho\rangle_{\textrm{TPM}}=\mathrm{D}_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\tilde{\rho}_{1})\neq\tilde{\rho}_{1}$ whereas $\langle\rho\rangle_{\textrm{DBN}}=\tilde{\rho}_{1}$. The Bayesian network approach therefore does not perturb the state of the working substance on average. In this sense, it is minimally invasive, in sharp contrast to the two-point measurement protocol. This difference arises because, in the Bayesian network scheme, measurements are performed in the eigenbasis of the density operator, in which the state is trivially diagonal. Figure 2 displays the trace distance, $d(\langle\rho\rangle,\tilde{\rho}_{1})$ nie10 , between the averaged measured state and the unmeasured state, as a function of the thermalization rate $\lambda$. These results indicate that two-point measurement experiments become increasingly invasive as the amount of coherence generated during the engine cycle grows for smaller $\lambda$.
Measurement backaction in the two-point measurement framework can, in fact, become so strong that it completely alters the operation regime of the machine. This behavior is illustrated in Fig. 3 as a function of the thermalization rate and the transverse driving strength. Depending on the parameter regime, the TPM-measured machine can run as an accelerator, that uses work to enhance the natural heat flow between the baths ($\langle w\rangle<0$, $\langle q_{\mathrm{h}}\rangle>0$, $\langle q_{\mathrm{c}}\rangle<0$), or a heater, that uses work to heat up both baths ($\langle w\rangle<0$, $\langle q_{\mathrm{h}}\rangle<0$, $\langle q_{\mathrm{c}}\rangle<0$) sol20 , while the unmeasured (and DBN-measured) machine functions as an engine, that uses the heat flow between the baths to output work ($\langle w\rangle>0$, $\langle q_{\mathrm{h}}\rangle>0$, $\langle q_{\mathrm{c}}\rangle<0$). The difference between the two, which can be related to a modification of the first law by the TPM-scheme (Supplemental Material), is especially marked for small thermalization rates $\lambda$ and large transverse driving strengths $g$.
Violation of fluctuation bounds.
Probability distributions not only give access to the mean but also to fluctuations around the mean. Recently, ”universal” bounds on such fluctuations have been proposed for quasistatic quantum heat engines in the form agarwalla_2021_2 ; agarwalla_2021 ; das23 ; moh23 ; moh23a ; watanabe_2025
$\eta^{(2)}=\frac{\sigma_{w}^{2}}{\sigma_{q_{\mathrm{h}}}^{2}}\leq\eta_{C}^{2}=\left(1-\frac{T_{\mathrm{c}}}{T_{\mathrm{h}}}\right)^{2},$ (5)
where $\sigma^{2}_{x}$ denotes the variance of the random variable $x$. Such upper bounds on the work output fluctuations are important, since they determine the stability of a thermal machine, both in the classical seifert_2018 ; ryabov_2018 and in the quantum lutz_2021 ; fei22 domains: an engine is usually considered stable, if the work output fluctuations are smaller than the mean. The explicit expressions of the variances for the two measurement schemes may be found in the Supplemental Material.
Equation (5) was derived assuming full thermalization and in the absence of quantum coherence agarwalla_2021_2 ; agarwalla_2021 ; das23 ; moh23 ; moh23a ; watanabe_2025 . Figure 4 reveals that the bound is actually violated in the presence of quantum coherence, and this for both measurement schemes, for finite relaxation times. Coherence is seen to augment work output fluctuations. We stress that the value of the transverse driving strength is here one order of magnitude smaller than the frequencies to ensure that the engine runs in the quasistatic regime.
Conclusions.
Quantum signatures manifest not only in the mean values of thermodynamic variables but also in their fluctuations. It is therefore crucial to develop an effective method to determine the energy exchange statistics of quantum machines with minimal measurement backaction, to preserve coherent features. We have shown that the dynamic Bayesian network approach is minimally invasive, in the sense that the averaged measured state exactly coincides with the unmeasured one, unlike the two-point measurement protocol. The measurement backaction of the latter can be so strong that it generally fails to reproduce the average work output of a coherent engine and may even dramatically alter its operation mode. In such cases, the measured motor may cease to function as a heat engine altogether. Furthermore, we have found that recently derived ”universal” fluctuation bounds do not necessarily apply to coherent engines. Our results are not restricted to the quantum Otto cycle, and hence offer a versatile platform for both theoretical and experimental investigations of the energy exchange statistics of coherent machines.
Acknowledgements. Financial support from the German Research Foundation DFG (Grant FOR 2724) and the Georg H. Endress Foundation is acknowledged.
Supplemental Material: Minimal-backaction work statistics of coherent engines
I Work and heat probability distributions: mean values and variances
We begin by explaining how to compute the work and heat probability distributions using the two-point measurement (TPM) scheme and the dynamic Bayesian network (DBN) approach. Their general form is den20s ; jia21s ; den21s ; lutz_2021s ; fei22s ; serra_2024s
$P(w)=\sum_{j,l,m,n}\delta[w+(e_{\mathrm{k}}^{l}-e_{\mathrm{e}}^{j})+(e_{\mathrm{e}}^{n}-e_{\mathrm{k}}^{m})]\,p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n}),$ (S1)
for the work probability distribution, and
$P(q_{\mathrm{h}})=\sum_{l,m}\delta[q_{\mathrm{h}}-(e_{\mathrm{k}}^{m}-e_{\mathrm{k}}^{l})]\,p_{q_{\mathrm{h}}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m}),$ (S2)
and
$P(q_{\mathrm{c}})=\sum_{n,r}\delta[q_{\mathrm{c}}-(e_{\mathrm{e}}^{r}-e_{\mathrm{e}}^{n})]\,p_{q_{\mathrm{c}}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}),$ (S3)
for the heat exchanged probability distributions with the hot and cold baths, respectively. Above, $e_{\mathrm{k},\mathrm{e}}^{j}$ is the $j$-th eigenvalue of the Hamiltonian $H_{\mathrm{k},\mathrm{e}}(t_{\mathrm{k},\mathrm{e}})$ and $p_{w}$, $p_{q_{\mathrm{h}}}$ and $p_{q_{\mathrm{c}}}$ are the joint probabilities of obtaining the corresponding eigenvalues. We note that $p_{w}$, $p_{q_{\mathrm{h}}}$ and $p_{q_{\mathrm{c}}}$ can all be obtained from a single probability distribution $p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ by computing the associated marginals:
$p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})$ $=\sum_{r}p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}),$ (S4)
$p_{q_{\mathrm{h}}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m})$ $=\sum_{r,n,j}p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}),$
$p_{q_{\mathrm{c}}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ $=\sum_{j,l,m}p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}).$
Operationally, $p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ represents the probability of sequentially obtaining the string of eigenvalues $\{e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}\}$ in an experiment in which the engine is set to run until its working substance reaches the steady state, then it is first measured after the cold isotherm, isentropically compressed, measured again, coupled to the hot bath, measured, isentropically expanded, measured, coupled to the cold bath and measured one last time. The working substance is then left alone to evolve until it reaches a steady state again and the experiment is repeated in order to obtain the probability of a different string of eigenvalues. The reason we compute $p_{w}$, $p_{q_{\mathrm{h}}}$ and $p_{q_{\mathrm{c}}}$ from the single probability distribution $p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ is so that the statistics of the work extracted are correlated with the ones of the heat exchanged, thus making a more realistic model than if all of them were computed individually and, therefore, completely uncorrelated.
Once the probability distributions are obtained, the mean values of their respective variables follow as
$\langle w\rangle=\int_{\mathbb{R}}dw\,w\,P(w)=-\sum_{j,l,m,n}[(e_{\mathrm{k}}^{l}-e_{\mathrm{e}}^{j})+(e_{\mathrm{e}}^{n}-e_{\mathrm{k}}^{m})]\,p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n}),$ (S5)
for the work, as well as
$\langle q_{\mathrm{h}}\rangle=\int_{\mathbb{R}}dq_{\mathrm{h}}\,q_{\mathrm{h}}\,P(q_{\mathrm{h}})=\sum_{l,m}(e_{\mathrm{k}}^{m}-e_{\mathrm{k}}^{l})\,p_{q_{\mathrm{h}}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m}),$ (S6)
and
$\langle q_{\mathrm{c}}\rangle=\int_{\mathbb{R}}dq_{\mathrm{c}}\,q_{\mathrm{c}}\,P(q_{\mathrm{c}})=\sum_{n,r}(e_{\mathrm{e}}^{r}-e_{\mathrm{e}}^{n})\,p_{q_{\mathrm{c}}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}),$ (S7)
for the exchanged heats. On the other hand, the corresponding variances are given by
$\sigma_{w}^{2}=\int_{\mathbb{R}}dw\,(w-\langle w\rangle)^{2}\,P(w)=\sum_{j,l,m,n}[(e_{\mathrm{k}}^{l}-e_{\mathrm{e}}^{j})+(e_{\mathrm{e}}^{n}-e_{\mathrm{k}}^{m})+\langle w\rangle]^{2}\,p_{w}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n}),$ (S8)
$\sigma_{q_{\mathrm{h}}}^{2}=\int_{\mathbb{R}}dq_{\mathrm{h}}\,(q_{\mathrm{h}}-\langle q_{\mathrm{h}}\rangle)^{2}\,P(q_{\mathrm{h}})=\sum_{l,m}[(e_{\mathrm{k}}^{m}-e_{\mathrm{k}}^{l})-\langle q_{\mathrm{h}}\rangle]^{2}\,p_{q_{\mathrm{h}}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m}),$ (S9)
and
$\sigma_{q_{\mathrm{c}}}^{2}=\int_{\mathbb{R}}dq_{\mathrm{c}}\,(q_{\mathrm{c}}-\langle q_{\mathrm{c}}\rangle)^{2}\,P(q_{\mathrm{c}})=\sum_{n,r}[(e_{\mathrm{e}}^{r}-e_{\mathrm{e}}^{n})-\langle q_{\mathrm{c}}\rangle]^{2}\,p_{q_{\mathrm{c}}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r}).$ (S10)
I.1 Two-point measurement scheme (TPM)
To evaluate the joint distribution $p(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ with the TPM scheme mukamel_2009s ; talkner_2011s , one starts with a single copy of the working substance and makes five consecutive measurements (interspaced by the four engine cycle branches) in the eigenbasis of the corresponding Hamiltonian at that point in time. We find
$p^{\mathrm{TPM}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})=\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\Pi_{\mathrm{e}}^{n}\,\mathcal{U}_{\mathrm{e}}\left\{\Pi_{\mathrm{k}}^{m}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\Pi_{\mathrm{k}}^{l}\,\mathcal{U}_{\mathrm{k}}\left(\Pi_{\mathrm{e}}^{j}\tilde{\rho}_{1}\Pi_{\mathrm{e}}^{j}\right)\Pi_{\mathrm{k}}^{l}\right]\Pi_{\mathrm{k}}^{m}\right\}\Pi_{\mathrm{e}}^{n}\right)\right].$ (S11)
The associated marginals are therefore
$p_{w}^{\mathrm{TPM}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})$ $=\text{tr}\left(\Pi_{\mathrm{e}}^{n}\,\mathcal{U}_{\mathrm{e}}\left\{\Pi_{\mathrm{k}}^{m}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\Pi_{\mathrm{k}}^{l}\,\mathcal{U}_{\mathrm{k}}\left(\Pi_{\mathrm{e}}^{j}\tilde{\rho}_{1}\Pi_{\mathrm{e}}^{j}\right)\Pi_{\mathrm{k}}^{l}\right]\Pi_{\mathrm{k}}^{m}\right\}\right),$ (S12)
$p_{q_{\mathrm{h}}}^{\mathrm{TPM}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m})$ $=\text{tr}\left\{\Pi_{\mathrm{k}}^{m}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\Pi_{\mathrm{k}}^{l}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)\Pi_{\mathrm{k}}^{l}\right]\right\},$
$p_{q_{\mathrm{c}}}^{\mathrm{TPM}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\Pi_{\mathrm{e}}^{n}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)\Pi_{\mathrm{e}}^{n}\right)\right].$
The mean values for work and heat are accordingly given by
$\langle w\rangle_{\mathrm{TPM}}$ $=\text{tr}[H_{\mathrm{k}}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{k}}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]+\text{tr}(H_{\mathrm{e}}\tilde{\rho_{1}})-\text{tr}[H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)],$ (S13)
$\langle q_{{\mathrm{h}}}\rangle_{\mathrm{TPM}}$ $=\text{tr}[H_{\mathrm{k}}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{k}}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)],$
$\langle q_{{\mathrm{c}}}\rangle_{\mathrm{TPM}}$ $=\text{tr}[H_{\mathrm{e}}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)],$
where we have used $\mathcal{T}_{t_{\mathrm{h,c}}}^{\mathrm{h,c}},\mathrm{D}_{\mathrm{k,e}}]=0$ and $\mathrm{D}_{\mathrm{k,c}}(H_{\mathrm{k,c}})=H_{\mathrm{k,c}}$. Interestingly, the first law for TPM-measured engine is
$\langle w\rangle_{\mathrm{TPM}}=\langle q_{{\mathrm{h}}}\rangle_{\mathrm{TPM}}+\langle q_{{\mathrm{c}}}\rangle_{\mathrm{TPM}}+\text{tr}\left\{H_{\mathrm{e}}\left[\tilde{\rho_{1}}-\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)\right]\right\},$ (S14)
where the last term measures the average deviation of the energy of the state of the working substance after a round of projective energy measurements at each point of the cycle. It is thus different from the unmeasured engine.
Moreover, the respective variances read
$\sigma_{w}^{2}$ $=\sigma_{q_{\mathrm{h}}}^{2}+\sigma_{q_{\mathrm{c}}}^{2}+\left\{\text{tr}\left\{H_{\mathrm{e}}^{2}\left[\tilde{\rho_{1}}-\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)\right]\right\}-\text{tr}\left\{H_{\mathrm{e}}\left[\tilde{\rho_{1}}-\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)\right]\right\}^{2}\right\}$ (S15)
$+\text{tr}[\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(H_{\mathrm{e}}^{2}\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{e}}^{2}\tilde{\rho}_{1}]-2\langle q_{{\mathrm{h}}}\rangle\langle q_{{\mathrm{c}}}\rangle-2(\langle q_{{\mathrm{h}}}\rangle+\langle q_{{\mathrm{c}}}\rangle)\text{tr}\left\{H_{\mathrm{e}}\left[\tilde{\rho_{1}}-\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)\right]\right\}$
$+2\,\text{tr}\{H_{\mathrm{e}}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[H_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\}+2\,\text{tr}\{[H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[H_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\}+2\,\text{tr}[H_{\mathrm{k}}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(H_{\mathrm{e}}\tilde{\rho}_{1}\right)]$
$-2\,\text{tr}[H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(H_{\mathrm{e}}\tilde{\rho}_{1}\right)]-2\,\text{tr}\{H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}[H_{\mathrm{e}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\}-2\,\text{tr}[H_{\mathrm{k}}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(H_{\mathrm{e}}\tilde{\rho}_{1}\right)]$
$\sigma_{q_{\mathrm{h}}}^{2}$ $=\text{tr}[H_{\mathrm{k}}^{2}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{k}}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]^{2}+\text{tr}[H_{\mathrm{k}}^{2}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{k}}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]^{2}$ (S16)
$+2\,\text{tr}[H_{\mathrm{k}}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\text{tr}[H_{\mathrm{k}}\,\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-2\,\text{tr}\{H_{\mathrm{k}}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[H_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\},$
$\sigma_{q_{\mathrm{c}}}^{2}$ $=\text{tr}[H_{\mathrm{e}}^{2}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{e}}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]^{2}$ (S17)
$+\text{tr}[H_{\mathrm{e}}^{2}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-\text{tr}[H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]^{2}$
$+2\,\text{tr}[H_{\mathrm{e}}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\text{tr}[H_{\mathrm{e}}\,\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]-2\,\text{tr}\{H_{\mathrm{e}}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[H_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}\left(\tilde{\rho}_{1}\right)]\}$
I.2 Dynamic Bayesian network approach (DBN)
In the case of the DBN protocol lutz_2020s ; par20s ; mic20as ; str20s ; zha22s ; lutz_2023s ; lutz_2024s , it is convenient to consider five copies of the working substance lutz_2023s : $\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}\otimes\tilde{\rho}_{1}$ (If only $p_{w}$ and $p_{q_{\mathrm{h}}}$ are required, only four copies are necessary). Given that the quantum engine reaches a steady state after a few cycles, in order to obtain the five copies, it is enough to initialize the engine with five arbitrary states $\rho$ and wait until they all reach the steady state $\tilde{\rho}_{1}$. Let $\tilde{\rho}_{i}=\sum_{\alpha}\lambda_{i}^{\alpha}\mathrm{P}_{i}^{\alpha}$ be the spectral decomposition of the state of the working substance at corner $i$ of the engine cycle. The measurement protocol to obtain $p^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})$ proceeds as follows:
  1. 1.
    The first two copies are (locally) projectively measured in the eigenbasis of $\tilde{\rho}_{1}$ and the last four copies are left to evolve. The unnormalized state after this step is:
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\otimes\tilde{\rho}_{2}\otimes\tilde{\rho}_{2}\otimes\tilde{\rho}_{2}.$ (S18)
  2. 2.
    The second and third copies are measured in the eigenbasis of $\tilde{\rho}_{2}$, and the last three copies are left to evolve. The updated state now is:
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\otimes\tilde{\rho}_{3}\otimes\tilde{\rho}_{3}.$ (S19)
  3. 3.
    The third and fourth copies are measured in the eigenbasis of $\tilde{\rho}_{3}$ and the last two copies are left to evolve to obtain:
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathrm{P}_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\mathrm{P}_{3}^{\gamma}\otimes\mathcal{U}_{\mathrm{e}}[\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}/\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})]\otimes\tilde{\rho}_{4}.$ (S20)
  4. 4.
    Copies $4$ and $5$ are measured in the eigenbasis of $\tilde{\rho}_{4}$ and the last copy is left to evolve. The state is updated to:
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathrm{P}_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\mathrm{P}_{3}^{\gamma}$ (S21)
    $\otimes\mathrm{P}_{4}^{\delta}\mathcal{U}_{\mathrm{e}}[\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}/\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})]\mathrm{P}_{4}^{\delta}\otimes\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}/\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})].$
  5. 5.
    A projective measurement is performed on the last copy in the eigenbasis of $\tilde{\rho}_{1}$ to complete the cycle and get the final state:
    $\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}\otimes\mathrm{P}_{2}^{\beta}\mathcal{U}_{\mathrm{k}}[\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}/\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})]\mathrm{P}_{2}^{\beta}\otimes\mathrm{P}_{3}^{\gamma}\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}[\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}/\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})]\mathrm{P}_{3}^{\gamma}$ (S22)
    $\otimes\mathrm{P}_{4}^{\delta}\mathcal{U}_{\mathrm{e}}[\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}/\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})]\mathrm{P}_{4}^{\delta}\otimes\mathrm{P}_{1}^{\epsilon}\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}[\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}/\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})]\mathrm{P}_{1}^{\epsilon}.$
The probability of the string of measurement outcomes $\{\lambda_{1}^{\alpha},\lambda_{2}^{\beta},\lambda_{3}^{\gamma},\lambda_{4}^{\delta},\lambda_{1}^{\epsilon}\}$ is then obtained by tracing over the un-normalized state:
$p(\lambda_{1}^{\alpha},\lambda_{2}^{\beta},\lambda_{3}^{\gamma},\lambda_{4}^{\delta},\lambda_{1}^{\epsilon})$ $=\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}$ (S23)
$\times\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left[\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\right\}.$
Additionally, the inferred probabilities of the energy eigenvalues given a partircular measurement outcome are
$p(e_{\mathrm{e}}^{j}|\lambda_{1}^{\alpha})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{j}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right],$ (S24)
$p(e_{\mathrm{k}}^{l}|\lambda_{2}^{\beta})$ $=\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right],$
$p(e_{\mathrm{k}}^{m}|\lambda_{3}^{\gamma})$ $=\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right],$
$p(e_{\mathrm{e}}^{n}|\lambda_{4}^{\delta})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right],$
$p(e_{\mathrm{e}}^{r}|\lambda_{1}^{\epsilon})$ $=\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
As a result, the probability distribution for the energy eigenvalues is
$p^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})=\sum_{\alpha,\beta,\gamma,\delta,\epsilon}p(\lambda_{1}^{\alpha},\lambda_{2}^{\beta},\lambda_{3}^{\gamma},\lambda_{4}^{\delta},\lambda_{1}^{\epsilon})\,p(e_{\mathrm{e}}^{j}|\lambda_{1}^{\alpha})\,p(e_{\mathrm{k}}^{l}|\lambda_{2}^{\beta})\,p(e_{\mathrm{k}}^{m}|\lambda_{3}^{\gamma})\,p(e_{\mathrm{e}}^{n}|\lambda_{4}^{\delta})\,p(e_{\mathrm{e}}^{r}|\lambda_{1}^{\epsilon}).$ (S25)
In this case, the marginals are
$p_{w}^{\mathrm{DBN}}(e_{\mathrm{e}}^{j},e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m},e_{\mathrm{e}}^{n})=$ $\sum_{\alpha,\beta,\gamma,\delta}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\},$ (S26)
$\times\text{tr}\left[\Pi_{\mathrm{e}}^{j}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right],$
$p_{q_{\mathrm{h}}}^{\mathrm{DBN}}(e_{\mathrm{k}}^{l},e_{\mathrm{k}}^{m})=$ $\sum_{\beta,\gamma}\text{tr}\left[\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}\right)\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{l}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[\Pi_{\mathrm{k}}^{m}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right],$
$p_{q_{\mathrm{c}}}^{\mathrm{DBN}}(e_{\mathrm{e}}^{n},e_{\mathrm{e}}^{r})=$ $\sum_{\delta,\epsilon}\text{tr}\left[\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}\right)\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{n}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[\Pi_{\mathrm{e}}^{r}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
The mean values of work and heat follow as
$\langle w\rangle_{{\mathrm{DBN}}}$ $=\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{3})-\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{2})+\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{1})-\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{4})=W$ (S27)
$\langle q_{\mathrm{h}}\rangle_{{\mathrm{DBN}}}$ $=\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{3})-\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{2})=Q_{\mathrm{h}}$
$\langle q_{\mathrm{c}}\rangle_{{\mathrm{DBN}}}$ $=\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{1})-\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{4})=Q_{\mathrm{c}}.$
The variances furthermore read
$\sigma_{w}^{2}$ $=\sigma_{q_{\mathrm{h}}}^{2}+\sigma_{q_{\mathrm{c}}}^{2}-2\langle q_{\mathrm{h}}\rangle\langle q_{\mathrm{c}}\rangle+2\sum_{\delta,\epsilon}\text{tr}\left[\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}\right)\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right]$ (S28)
$-2\sum_{\alpha,\beta,\gamma,\delta}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}$
$\times\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]$
$+2\sum_{\beta,\gamma,\delta}\text{tr}\left[\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}\right)\right]\,\text{tr}\left\{\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left[\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\right\}\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]$
$+2\sum_{\alpha,\beta,\gamma}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left\{\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left[\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\right\}\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]$
$-2\sum_{\alpha,\beta}\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})\,\text{tr}\left\{\mathrm{P}_{2}^{\beta}\,\mathcal{U}_{\mathrm{k}}\left[\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\right\}\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1}\mathrm{P}_{1}^{\alpha}}{\text{tr}(\mathrm{P}_{1}^{\alpha}\tilde{\rho}_{1})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]$
$-2\sum_{\gamma,\delta}\text{tr}\left[\mathrm{P}_{4}^{\delta}\,\mathcal{U}_{\mathrm{e}}\left(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}\right)\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right].$
$\sigma_{q_{\mathrm{h}}}^{2}$ $=\text{tr}(H_{\mathrm{k}}^{2}\tilde{\rho}_{3})-\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{3})^{2}+\text{tr}(H_{\mathrm{k}}^{2}\tilde{\rho}_{2})-\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{2})^{2}$ (S29)
$+2\,\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{3})\text{tr}(H_{\mathrm{k}}\tilde{\rho}_{2})-2\sum_{\beta,\gamma}\text{tr}\left[\mathrm{P}_{3}^{\gamma}\,\mathcal{T}_{t_{\mathrm{h}}}^{\mathrm{h}}\left(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}\right)\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2}\mathrm{P}_{2}^{\beta}}{\text{tr}(\mathrm{P}_{2}^{\beta}\tilde{\rho}_{2})}\right]\,\text{tr}\left[H_{\mathrm{k}}\,\frac{\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3}\mathrm{P}_{3}^{\gamma}}{\text{tr}(\mathrm{P}_{3}^{\gamma}\tilde{\rho}_{3})}\right].$
$\sigma_{q_{\mathrm{c}}}^{2}$ $=\text{tr}(H_{\mathrm{e}}^{2}\tilde{\rho}_{1})-\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{1})^{2}+\text{tr}(H_{\mathrm{e}}^{2}\tilde{\rho}_{4})-\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{4})^{2}$ (S30)
$+2\,\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{1})\text{tr}(H_{\mathrm{e}}\tilde{\rho}_{4})-2\sum_{\delta,\epsilon}\text{tr}\left[\mathrm{P}_{1}^{\epsilon}\,\mathcal{T}_{t_{\mathrm{c}}}^{\mathrm{c}}\left(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}\right)\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4}\mathrm{P}_{4}^{\delta}}{\text{tr}(\mathrm{P}_{4}^{\delta}\tilde{\rho}_{4})}\right]\,\text{tr}\left[H_{\mathrm{e}}\,\frac{\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1}\mathrm{P}_{1}^{\epsilon}}{\text{tr}(\mathrm{P}_{1}^{\epsilon}\tilde{\rho}_{1})}\right].$
II Equality between the TPM and DBN probability distributions
In this section, we will show that the conditions
$\begin{cases}&\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\rho)=\mathrm{D}_{\mathrm{k}}\mathcal{U}_{\mathrm{k}}\mathrm{D}_{\mathrm{e}}(\rho)\quad\text{or}\quad t_{\mathrm{h}}\to\infty\\ &\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\rho)=\mathrm{D}_{\mathrm{e}}\mathcal{U}_{\mathrm{e}}\mathrm{D}_{\mathrm{k}}(\rho)\quad\text{or}\quad t_{\mathrm{c}}\to\infty.\end{cases}$ (S31)
imply that the state of the working substance is diagonal in the basis of the corresponding Hamiltonian at each corner of the cycle and, therefore, that the probability distributions computed with the TPM or DBN schemes are the same. In order to do so, first note that generically the generalized amplitude damping map (as a function of the thermalization rate instead of the time) can be written as srikanth_2024
$\mathcal{T}_{\lambda}(\rho)=(1-\lambda)\rho+\lambda\,\sigma-2\sqrt{1-\lambda}(1-\sqrt{1-\lambda})\sum_{j=1}^{d}p_{j}\mathcal{D}_{\Pi^{j}}(\rho),$ (S32)
where $\lambda\in[0,1]$, $\sigma=\sum_{j}p_{j}\Pi^{j}$ is the unique fixed point of $\mathcal{T}_{\lambda}$, and $\mathcal{D}_{\Pi^{j}}(\cdot)=\Pi^{j}\cdot\Pi^{j}-\{\Pi^{j},\cdot\}/2$ is the usual Lindblad dissipator. In the case of the map describing the interaction with the cold reservoir, $\mathcal{T}_{\lambda}^{\mathrm{c}}$, the fixed point is $\sigma=\text{exp}[-H_{\mathrm{e}}(t_{\mathrm{e}})/T_{\mathrm{c}}]/\text{tr}\{\text{exp}[-H_{\mathrm{e}}(t_{\mathrm{e}})/T_{\mathrm{c}}]\}$, and the projectors are $\{\Pi_{\mathrm{e}}^{j}\}_{j}$. Analogously, for the interaction with the hot reservoir, $\mathcal{T}_{\lambda}^{\mathrm{h}}$, $\sigma=\text{exp}[-H_{\mathrm{k}}(t_{\mathrm{k}})/T_{\mathrm{h}}]/\text{tr}\{\text{exp}[-H_{\mathrm{k}}(t_{\mathrm{k}})/T_{\mathrm{h}}]\}$ and the projectors are $\{\Pi_{\mathrm{k}}^{j}\}_{j}$. It is straightforward to check that, if $\mathrm{D}(\cdot)=\sum_{j}\Pi^{j}\cdot\Pi^{j}$ is the dephasing map in the eigenbasis of the fixed point of $\mathcal{T}_{\lambda}$, then $\mathrm{D},\mathcal{T}_{\lambda}]=0$. Thus, diagonal states in that basis are preserved by $\mathcal{T}_{\lambda}$. On the other hand, if $\rho$ was not diagonal in that basis, then $\mathcal{T}_{\lambda}(\rho)$ is also not diagonal unless $\lambda=1$. Therfore, $\mathcal{T}_{\lambda}(\rho)$ is diagonal if $\rho$ was already diagonal or if $\lambda=1$, as it was stated in the main text. Now, assuming the conditions in Eq. (S31) hold, a direct computation shows that diagonal states in the basis of $H_{\mathrm{e}}(t_{\mathrm{e}})$ are preserved during the engine cycle, $\Lambda_{1}D_{\mathrm{e}}(\rho)=D_{\mathrm{e}}\Lambda_{1}D_{\mathrm{e}}(\rho)$, which immediately leads to $\Lambda_{1}^{n}D_{\mathrm{e}}(\rho)=D_{\mathrm{e}}\Lambda_{1}^{n}D_{\mathrm{e}}(\rho)$. By taking the limit $n\to\infty$ and using the fact that the fixed point of $\Lambda_{1}$ is unique, the result is that the steady state $\tilde{\rho}_{1}$ is diagonal in the eigenbasis of $H_{\mathrm{e}}(t_{\mathrm{e}})$, $\tilde{\rho}_{1}=D_{\mathrm{e}}(\tilde{\rho}_{1})$. By the first line of Eq. (S31), the state of the working substance is diagonal in the eigenbasis of $H_{\mathrm{k}}(t_{\mathrm{k}})$ after the isentropic compression and, therefore, after the hot isothermal isochore too. The second line in Eq. (S31) then implies that it is diagonal in the eigenbasis of $H_{\mathrm{e}}(t_{\mathrm{e}})$ after the isentropic expansion. Consequently, the state of the working substance is diagonal in the corresponding Hamiltonian basis in the four corners of the cycle and, as a result, the TPM and DBN protocols are equivalent.
In Fig. S1 we show three generic examples of the work probability distributions $P_{\text{TPM}}$ and $P_{\text{DBN}}$. Note they coincide if $g=0$ or if $t_{\mathrm{h},\mathrm{c}}\to\infty$.
Supplementary Figure S1: Work probability distributions $P(w)$ of the engine cycle computed using TPM (red) and DBN (blue), with their respective means, $\langle w\rangle_{\mathrm{TPM}}$ and $\langle w\rangle_{\mathrm{DBN}}$, compared to the unmeasured value of work $W$. a) In the generic case, where coherences in the energy basis are dynamically induced, $g=9$, and the working substance does not fully thermalize, $t_{\mathrm{h}}=t_{\mathrm{c}}=0.92$, both probability distributions are different and only the one computed with DBN reproduces the correct mean value for the work. b) In the absence of dynamically induced coherences, $g=0$ and $t_{\mathrm{h}}=t_{\mathrm{c}}=0.92$, or c) if the working substance thermalizes, $g=9$ and $t_{\mathrm{h},\mathrm{c}}\to\infty$, TPM and DBN are equivalent. Parameters are $\omega_{\mathrm{h}}=10$, $T_{\mathrm{h}}=14$, $\omega_{\mathrm{c}}=0.5$, $T_{\mathrm{c}}=0.1$, $g=9$, $t_{\mathrm{h}}=t_{\mathrm{c}}=0.92$.
III Comparison between the TPM and DBN protocols
III.1 Kullback-Leibler divergence and relative entropy of coherence
[…]
Supplementary Figure S2: a) Kullback-Leibler divergence, $D_{\textrm{KL}}(P_{\textrm{DBN}}||P_{\textrm{TPM}})$, between the DBN and the TPM work probability distributions. b) Relative entropy of coherence $\mathcal{C}$ of the steady state before the isentropic compression step, $\tilde{\rho}_{1}$. c) Relative entropy of coherence $\mathcal{C}$ of the steady state before the isentropic expansion step, $\tilde{\rho}_{3}$. The regions in which the two probability distributions differ the most (a) coincide with the zones where the steady state has the most coherence (bc). Parameters are $\omega_{\mathrm{h}}=10$, $T_{\mathrm{h}}=14$, $\omega_{\mathrm{c}}=0.5$, $T_{\mathrm{c}}=0.1$, and $\lambda_{\mathrm{c}}=\lambda_{\mathrm{h}}=\lambda$.
III.2 Backaction and trace distance between measured and unmeasured steady states
In the main text, we further showed that the DBN scheme is minimally invasive by evaluating the trace distance, $d(\langle\rho\rangle,\tilde{\rho}_{1})$, between the averaged measured state $\langle\rho\rangle$ and the unmeasured steady state $\tilde{\rho}_{1}$. This trace distance vanishes for the DBN protocol but not for the TPM scheme, as seen in Fig. 2 as a function of the thermalization rate $\lambda$, but again for a fixed value of the transverse driving strength $g$. We here complement these findings with the contour plot of $d_{\text{TPM}}(\langle\rho\rangle,\tilde{\rho}_{1})$ as $g$ is varied (Fig. S3).
Supplementary Figure S3: Trace distance $d(\langle\rho\rangle,\tilde{\rho}_{1})$ between the average state $\langle\rho\rangle$ obtained after performing a round of successive TPM measurements starting and ending with the state after the cold isochore and the steady state $\tilde{\rho}_{1}$ of the unmeasured engine. Parameters are $\omega_{\mathrm{h}}=10$, $T_{\mathrm{h}}=14$, $\omega_{\mathrm{c}}=0.5$, $T_{\mathrm{c}}=0.1$, $g=9$, and $\lambda_{\mathrm{c}}=\lambda_{\mathrm{h}}=\lambda$.
IV Violation of ”universal” fluctuation bounds
Finally, we found that ”universal” fluctuation bounds put forward for quasistatic engines agarwalla_2021_2s ; agarwalla_2021s ; das23s ; moh23s ; moh23as ; watanabe_2025s can actually be violated in coherent engines (Fig. 4 of the main text), for both measurement protocols. We complement this figure by additional contour plots for the ratio $\eta^{(2)}/\eta_{C}^{2}$, for varying transverse field strength $g$ (Fig. S4). Figure S5 shows that these violations are related to the presence of coherence in the cycle (even though these are small in the quasistatic regime).
Supplementary Figure S4: Quotient $\eta^{(2)}/\eta_{C}^{2}$ evaluated with a) the DBN protocol and with b) the TPM scheme, in the quasistatic regime. The white dashed lines indicate the set of values such that $\eta^{(2)}/\eta_{C}^{2}=1$. Every value above that line adheres to the bound, while everything below it violates it. Parameters used are $\omega_{\mathrm{h}}=1$, $T_{\mathrm{h}}=1.2$, $\omega_{\mathrm{c}}=0.85$, $T_{\mathrm{c}}=1$, and $\lambda_{\mathrm{c}}=\lambda_{\mathrm{h}}=\lambda$.
(a) Relative entropy of coherence $\mathcal{C}$ of the steady state before the isentropic compression step, $\tilde{\rho}_{1}$.
(b) Relative entropy of coherence $\mathcal{C}$ of the steady state before the isentropic expansion step, $\tilde{\rho}_{3}$.
Supplementary Figure S5: a) Relative entropy of coherence $\mathcal{C}$ of the steady state before the isentropic compression step, $\tilde{\rho}_{1}$, and b) before the isentropic expansion step, $\tilde{\rho}_{3}$, in the quasistatic regime. Same parameters as in Fig. S4.